# Efficient Backdoor Attacks for Deep Neural Networks in Real-world Scenarios

### v1

@@ Abstract
Recent deep neural networks (DNNs) have come to rely on vast amounts of training data, providing an opportunity for malicious attackers to exploit and contaminate the data to carry out backdoor attacks. These attacks significantly undermine the reliability of DNNs. However, existing backdoor attack methods make unrealistic assumptions, assuming that all training data comes from a single source and that attackers have full access to the training data. In this paper, we address this limitation by introducing a more realistic attack scenario where victims collect data from multiple sources, and attackers cannot access the complete training data. We refer to this scenario as data-constrained backdoor attacks. In such cases, previous attack methods suffer from severe efficiency degradation due to the entanglement between benign and poisoning features during the backdoor injection process.
To tackle this problem, we propose a novel approach that leverages the pre-trained Contrastive Language-Image Pre-Training (CLIP) model. We introduce three CLIP-based technologies from two distinct streams: Clean Feature Suppression, which aims to suppress the influence of clean features to enhance the prominence of poisoning features, and Poisoning Feature Augmentation, which focuses on augmenting the presence and impact of poisoning features to effectively manipulate the model’s behavior.
To evaluate the effectiveness, harmlessness to benign accuracy, and stealthiness of our method, we conduct extensive experiments on 3 target models, 3 datasets, and over 15 different settings. The results demonstrate remarkable improvements, with some settings achieving over 100% improvement compared to existing attacks in data-constrained scenarios. Our research contributes to addressing the limitations of existing methods and provides a practical and effective solution for data-constrained backdoor attacks.
11footnotetext: Equal Contribution.
@@ 1 Introduction
Deep neural networks (DNNs) are widely utilized and powerful machine learning algorithms inspired by the structure and functioning of the human brain. They excel at learning intricate patterns in data, making them invaluable for various applications such as image recognition [21, 17], natural language processing [33, 68], image generation [20, 30], and anomaly detection [64, 45]. However, the effectiveness of DNNs heavily relies on the quantity and quality of the training data. For instance, Stable Diffusion [49], a generative model with 983 million parameters, owes its success in image generation tasks to pre-training on 5 billion image-text pairs. Similarly, GPT-3 [3], a language model with 175 billion parameters, owes its efficacy in diverse language processing tasks to pre-training on an extensive corpus of 45 TB of text data. As the demand for data continues to rise, many users and businesses resort to third-party sources or online collections as a convenient means of acquiring the necessary data. However, recent studies [43, 18, 2, 25, 64, 15] have demonstrated that such practices can be maliciously exploited by attackers to contaminate the training data, significantly impairing the functionality and reliability of trained models.
The growing adoption of neural networks across different domains has made them an attractive target for malicious attacks. One particular attack technique gaining attention is the backdoor attack [15, 24, 26, 9, 27, 41, 2]. In backdoor attacks, a neural network is deliberately injected with a hidden trigger by introducing a small number of poisoning samples into the benign training set during the training. Once the model is deployed, the attacker can activate the backdoor by providing specific inputs containing the hidden trigger, causing the model to produce incorrect results. Backdoor attacks continue to present a significant and pervasive threat across multiple sectors, including image classification [16], natural language processing [43], speaker verification [67], malware detection [24], and video recognition [69]. In this paper, we focus on the widely studied field of image classification.
@@
However, it is important to note that all of previous backdoor attacks are based on an assumption that may be too broad in practice. They assume that all training data has been collected from a single source, and the collected source has been poisoned by attacker (As depicted in the O2O data collection mode in Fig. 1). In this case, the attacker can full access to the entire training data. While this assumption allows attackers to easily poison the training data, it does not accurately reflect real-world attack scenarios. To illustrate this point, let’s consider a scenario where victims possess a private dataset with limited samples. To compensate for the limited data, victims may need to augment their dataset by collecting additional data from multiple sources on the internet (referred to as the public dataset) and combine it with their private dataset to form the training set. As shown in the M2O data collection mode in Fig. 1, part of the sources may be poisoned by attackers, secretly***This scenario is common in the context of large models training. For instance, Stable Diffusion [49], a popular generative model, pre-trains on 5 billion image-text pairs. Similarly, GPT-3 [3], a language model with 175 billion parameters, pre-trains on an extensive corpus of 45 TB of text data. Both of them need to crawl large amounts of data from different websites to complete complex training tasks.. In this case, the attackers cannot access the private dataset and can only manipulate a portion of the public dataset to carry out the poisoning process. Consequently, a discrepancy arises between the distribution of the poisoning data and the distribution of the training data, which differs from the previous pipeline of poisoning attacks.
@@
In this paper, we address a more realistic backdoor attack scenario called data-constrained backdoor attacks, where the attackers do not have access to the entire training set. To be more precise, we classify data-constrained backdoor attacks into three types based on different types of data sources: number-constrained backdoor attacks, class-constrained backdoor attacks, and domain-constrained backdoor attacksThese three backdoor attacks are more realistic attack scenario, where the data provided by each data source is independently and identically distribution in number-constrained backdoor attacks, each data source provides data belonging to different categories in class-constrained backdoor attacks, and each data source provides data from different domains in domain-constrained backdoor attacks. More description can be found in Sec. 3.. Upon investigation, we have discovered that existing attack methods exhibit significant performance degradation when dealing with these data-constrained backdoor attacks. We propose that the entanglement between benign and poisoning features is a crucial factor contributing to this phenomenon. Entanglement refers to the neural networks utilizing both benign and poisoning features to make decisions for poisoning samples. However, this approach is not efficient for backdoor attacks. Ideally, an efficient backdoor attack should solely rely on the poison feature generated by the trigger to make decisions, irrespective of how the benign feature is expressed.
To address the aforementioned challenge and enhance the efficiency of poisoning attacks in data-constrained backdoor scenarios, we introduce two streams: Clean Feature Suppression and Poisoning Feature Augmentation. In the Clean Feature Suppression stream, we focus on reducing the influence of clean features during the poisoning process. On the other hand, in the Poisoning Feature Augmentation stream, we aim to amplify the expression of poisoning features. To achieve these goals, we propose three techniques utilizing the pre-trained Contrastive Language-Image Pre-Training (CLIP) [46] model.
i) CLIP-based Clean Feature Erasing (CLIP-CFE) is designed to suppress the expression of clean features. It leverages the capabilities of the CLIP model to identify and minimize the impact of clean features in the poisoning process.
ii) CLIP-based Universal Adversarial Perturbations (CLIP-UAP) focuses on poisoning feature augmentation. It employs the CLIP model to generate universal adversarial perturbations that effectively enhance the expression of poisoning features in the training data.
iii) CLIP-based Contrastive Feature Augmentation (CLIP-CFA) also falls under the poisoning feature augmentation stream. This technique utilizes the CLIP model to perform contrastive feature augmentation, enhancing the power of poisoning features and improving the effectiveness of the backdoor attack.
Our main contributions are summarized as follows.
  • We present a novel and contemporary backdoor attack scenario called data-constrained backdoor attacks to image classification. Data-constrained backdoor attacks assume that attackers lack access to the entire training data, making it a versatile and practical attack with broad applicability.
  • Through a systematic analysis of previous attack methods, we identify the entanglement between poisoning and benign features as the primary contributing factor to their performance degradation.
@@
  • To address this issue, we introduce the pre-trained CLIP model into the field of backdoor attacks for the first time. We propose three innovative technologies: CLIP-CFE, CLIP-UAP, and CLIP-CFA. Extensive evaluations conducted on 3 datasets and 3 target models, and over 15 different settings demonstrate the significant superiority of our proposed CLIP-UAP and CLIP-CFA over existing backdoor attacks. Furthermore, CLIP-CFE complements existing attack methods and can be seamlessly integrated with them, resulting in further efficiency improvements.
2 Background
Here we first summary the common pipeline of backdoor attacks on neural networks, and then we introduce the Contrastive Language-Image Pre-Training (CLIP) Model that has been adopted in our method.
@@ 2.1 Backdoor Attacks on Neural Networks
2.1.1 General Pipeline of Backdoor Attacks
@@
Backdoor injection. In this step, The attackers mix the poisoning set $\mathcal{P}$ into the clean training set $\mathcal{D}$ and release the new dataset. The victims download the poisoning dataset and use it to train their own DNN models [16]:
@@
Backdoor activation. In this step, the victims deploy their compromised DNN models on model-sharing platforms and model-selling platforms, such as Model Zoo and AWS Marketplace. The compromised model behaves normally when presented with benign inputs, but attackers can manipulate its predictions to align with their malicious objectives by providing specific samples containing pre-defined triggers.
2.1.2 Examples of Backdoor Attacks
Here, we present three popular backdoor attack methods that serve as the baseline for our preliminary experiments, providing insight into the motivation discussed in Sec. 3. All attacks follow the pipeline described in Sec. 2.1.1.
BadNets [16]. BadNets [16] is the pioneering backdoor attack in deep learning and is often used as a benchmark for subsequent research. It utilizes a $2\times 2$ attacker-specified pixel patch as the universal trigger pattern attached to benign samples.
Blended [5] Chen et al. [5] first discuss the requirement for invisibility in backdoor attacks. They propose that the poisoning image should be visually indistinguishable from its benign counterpart to evade human inspection. To meet this requirement, they introduce a blending strategy where poisoning images are created by blending the backdoor trigger with benign images. Formally, the poison generator can be formulated as $\mathcal{T}(x,t)=\lambda\cdot t+(1-\lambda)\cdot x$, where $\lambda$ represents the blend ratio (we set $\lambda=0.15$ for all experiments in this paper), and $t$ is an attacker-specified benign image serving as the universal trigger pattern.
Universal Adversarial Perturbations (UAP) [70] Inspired by Universal Adversarial Perturbations (UAPs) in adversarial examples, some studies [70, 26, 9] propose optimizing a UAP on a pre-trained clean model as the natural trigger, formulated as $\mathcal{T}(x,t)=x+t$, where $t$ is a pre-defined UAP serving as the universal trigger pattern. It’s worth noting that UAP-based backdoor attacks require a clean model pre-trained on the entire training set, which is not suitable for the discussed settings. However, to better explain our motivation that previous technologies exhibit significant performance degradation in data-constrained backdoor attacks, we assume the availability of a clean model pre-trained on the original training dataset in this section. It is important to acknowledge that this assumption does not hold in an actual attack scenario.
(a) Number-constrained Scenario
(b) Class-constrained Scenario
(c) Domain-constrained Scenario
2.2 Contrastive Language-Image Pre-Training (CLIP) Model
Our method introduces the Contrastive Language-Image Pre-Training (CLIP) [46] model into backdoor injection and we introduce it here. CLIP is a revolutionary deep learning model developed by OpenAI that is designed to connects texts and images by bringing them closer in a shared latent space, under a contrastive learning manner. The CLIP model is pre-trained on 400 million image-text pairs harvested from the Web, containing two encoder: CLIP text encoder ^$\hat{\mathcal{E}}_{t}(\cdot)$ and CLIP image encoder ^$\hat{\mathcal{E}}_{i}(\cdot)$. These encoders project the text and image to the CLIP common embedded feature space. Since natural language is able to express a much wider set of visual concepts, it contains ability to generalize across a wide range of tasks and domains, such as text-driven image manipulation [44], zero-shot classification [6], domain generalization [42]. To our best knowledge, our paper is the first study to explore the usage of CLIP model in the security community.
@@ 3 Data-constrained Backdoor Attacks
Here we first show the considered pipeline of data-constrained backdoor attacks, and than illustrate the performance degradation of previous attack methods on proposed data-constrained backdoor attacks. Finally, we attribute the degradation to the entanglement between the benign and poisoning features during the poisoning injection.
@@ 3.1 Preliminaries
Previous methods [5, 16, 34, 27, 60, 12, 41, 39, 59] have commonly followed the attack pipeline outlined in Sec. 2.1.1. However, this widely adopted pipeline relies on an overly loose assumption that all training data is collected from a single source and that the attacker has access to the entire training data, which is often not the case in real-world attack scenarios. In this paper, we focus on a more realistic scenario: Data-constrained Backdoor Attacks, which necessitates the definition of a modified attack pipeline as follows:
@@
Pipeline of data-constrained backdoor attacks. Similar to the general pipeline of backdoor attacks on neural networks, the proposed pipeline also consists of three steps: poisoning set generation, backdoor injection, and backdoor activation. The backdoor injection and activation steps remain unchanged from the previous pipeline. However, in the poisoning set generation step, data-constrained backdoor attacks differ from the previous pipeline. Instead of assuming access to the entire dataset $\mathcal{D}$, data-constrained attacks only assume access to a clean training set $\mathcal{D^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,N^{\prime}\}$, which follows a different data distribution from $\mathcal{D}$. To address this, the attacker randomly selects a subset $\mathcal{P^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,P\}$ from the accessible dataset $\mathcal{D^{\prime}}$, and creates the corresponding poisoning set $\mathcal{P}=\{(x^{\prime}_{i},k)|x^{\prime}_{i}=\mathcal{T}(x_{i},t),(x_{i},y_{i})\in\mathcal{P^{{}^{\prime}}},i=1,\cdots,P\}$. Additionally, based on the different constraints imposed by the accessible training set $\mathcal{D^{\prime}}$, data-constrained backdoor attacks are further categorized into three types: Number-constrained Backdoor Attacks, Class-constrained Backdoor Attacks, and Domain-constrained Backdoor Attacks. The details can be found in Sec. 3.2, Sec 3.3, and Sec. 3.4, respectively.
Experimental settings. To evaluate the performance of three backdoor attack methods (BadNets, Blended, and UAP) under data-constrained scenarios, we conduct experiments on the CIFAR-10 dataset. Specifically, we consider three types of data constraints: number, class, and domain. The settings for the poisoning attacks follow those described in Sec. 2.1.2. In all attacks, we set the attack-target label $k$ to category 0. For our experiments, we select the VGG-16 model as the victim model and employ SGD as the optimizer with a weight decay of 5e-4 and a momentum of 0.9. The batch size is set to 256, and the initial learning rate is set to 0.01. The learning rate is multiplied by 0.1 at the 35-th and 55-th epochs, and the training is conducted for a total of 70 epochs.
3.2 Number-constrained Backdoor Attacks
Definition. Let $\mathcal{D^{\prime}}$ denote the data manipulable by the malicious source, and $\mathcal{D}$ represent all the data available to the data collector. In the number-constrained scenario, as illustrated in Fig. 2 (a), the data collector gathers data from multiple sources, including both malicious and benign sources, to form $\mathcal{D}$. The data provided by each data source is independently and identically distributed. In other words, $\mathcal{D}$ and $\mathcal{D^{\prime}}$ belong to the same distribution, but in terms of quantity, $N^{\prime}<N$. The setting of number-constrained backdoor attacks is similar to that of data-efficient backdoor attacks discussed in previous studies [60, 70]. Both aim to improve the Attack Success Rate (ASR) under a low poisoning rate. However, previous studies assumed that the attacker has access to the entire training set $\mathcal{D}$, which enables efficient trigger design and sample selection. For example, some studies [70] draw inspiration from Universal Adversarial Perturbations (UAPs) in adversarial examples and propose to optimize a UAP on a clean model pre-trained on the training set as the natural trigger. Xia et al. [60] enhance the poisoning efficiency in backdoor attacks by selecting poisoning data from the entire training set. Although these methods have achieved remarkable results, they cannot be directly applied to number-constrained backdoor attacks due to the lack of access to the entire training set.
@@ 3.2 Number-constrained Backdoor Attacks
Experimental results. In this section, we investigate the performance degradation of previous studies in number-constrained backdoor attacks. As shown in Fig. 3, the attack success rate experiences a significant decrease as the number ($P$) of poisoning samples decreases, particularly for Blended backdoor attacks. It is worth noting that Universal Adversarial Perturbations (UAP) achieves relatively favorable results even with a low poisoning rate. This can be attributed to the utilization of a proxy model that is pre-trained on the entire training set ($\mathcal{D}$). However, in our settings, UAP is not accessible, and we present the results for UAP to effectively demonstrate the performance degradation even when a pre-trained proxy model is available.
Figure 3: The attack success rate (ASR) in the number-constrained backdoor attacks. The abscissa is the number ($P$) of samples in the poisoning set $\mathcal{P^{\prime}}$. The experiment is repeated 5 times, and the solid lines represent the mean results.
3.3 Class-constrained Backdoor Attacks
Definition. In the class-constrained scenario, let $\mathcal{D^{\prime}}$ represent the data manipulable by the malicious source, and $\mathcal{D}$ denote all the data available to the data collector. As depicted in part (b) of Fig. 2, the data collector gathers data from multiple sources, including both malicious and benign sources, to form $\mathcal{D}$. Each data source provides data belonging to different categories, resulting in $\mathcal{D^{\prime}}$ containing only a subset of categories present in $\mathcal{D}$. Therefore, $\mathcal{D}$ and $\mathcal{D^{\prime}}$ follow distinct distributions. More specifically, the accessible clean training set $\mathcal{D^{\prime}}\subset X\times Y^{\prime}(\mathcal{D^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,N^{\prime}\})$ is a subset of the entire training set $\mathcal{D}\subset X\times Y(\mathcal{D}=\{(x_{i},y_{i})|i=1,\cdots,N\})$, where $Y^{\prime}\subset Y=\{1,2,\cdots,C\}$. Class-constrained backdoor attacks can be seen as a general setting of clean-label backdoor attacks [54, 50, 53]. In clean-label backdoor attacks, the accessible clean training set $\mathcal{D^{\prime}}$ is defined as $\mathcal{D^{\prime}}\subset X\times Y^{\prime}$, where $Y^{\prime}=\{k\}$ and $k$ represents the attack-target label.
Experimental results. In this section, we explore the performance degeneration of previous studies in class-constrained backdoor attacks. As illustrated in Fig. 4, attack success rate decreases as the number of class ($C^{\prime}$) in the poisoning set decreases, which is the similar as experimental results on the number-constrained backdoor attacks.
@@ 3.3 Class-constrained Backdoor Attacks
Figure 4: The attack success rate (ASR) in the class-constrained backdoor attack. The experiments are conducted with triggers BadNets, Blended, and UAP, with a poisoning rate of $2\%$ ($P=1000$) for each. The x-axis represents the number of classes ($|Y^{\prime}|$) in the poisoning set $\mathcal{P^{\prime}}$. Specifically, ’1 (1)’ and ’1 (0)’ denote dirty-label single-class ($Y^{\prime}=\{c\},c\neq k$) and clean-label single-class ($Y^{\prime}=\{k\}$), respectively. The experiment is repeated 5 times, and the solid lines indicate the average results.
3.4 Domain-constrained Backdoor Attacks
Definition. In the domain-constrained scenario, as depicted in part (c) of Fig. 2, the data collector gathers data from multiple sources (both malicious and benign) to form $\mathcal{D}$. Each data source provides data from a different domain, resulting in $\mathcal{D^{\prime}}$ containing only a subset of the domains present in $\mathcal{D}$. Consequently, $\mathcal{D}$ and $\mathcal{D^{\prime}}$ belong to different distributions. We examine an extreme scenario in domain-constrained backdoor attacks, where the test dataset follows the same distribution as the benign source ($\mathcal{D}\setminus\mathcal{D^{\prime}}$) and is outside the domain of the malicious source $\mathcal{D^{\prime}}\subset X\times Y^{\prime}$ ($\mathcal{D^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,N^{\prime}\}$).
Figure 5: The attack success rate (ASR) in the domain-constrained backdoor attack. The poisoning rates of experiments with trigger BadNets, Blended, and UAP are $2\%$ ($P=1000$), $2\%$ ($P=1000$), and $1\%$ ($P=500$), respectively. The abscissa is the domain rate that represents the proportion of poisoning sets sampled from $\mathcal{D}\setminus\mathcal{D^{\prime}}$ and $\mathcal{D^{\prime}}$). The experiment is repeated 5 times, and the solid lines represent the mean runs.
Experimental results. To simulate the domain-constrained scenario, we conducted experiments with the following settings in this section: we designate the CIFAR-10 dataset as the benign source, the ImageNet dataset as the malicious source, and evaluated the attack performance on the CIFAR-10 dataset. Fig. 5 illustrates the results, showing a decrease in the attack success rate as the domain rate (the proportion of poisoning sets sampled from $\mathcal{D}\setminus\mathcal{D^{\prime}}$ and $\mathcal{D^{\prime}}$) in the poisoning set decreases. This observation aligns with the experimental findings in the number-constrained and class-constrained backdoor attacks.
(a) ASR of three backdoor attacks under the same poisoning rate ($p=1000$) on the CIFAR-10 dataset.

### v2

@@ Abstract
Recent deep neural networks (DNNs) have came to rely on vast amounts of training data, providing an opportunity for malicious attackers to exploit and contaminate the data to carry out backdoor attacks. However, existing backdoor attack methods make unrealistic assumptions, assuming that all training data comes from a single source and that attackers have full access to the training data. In this paper, we introduce a more realistic attack scenario where victims collect data from multiple sources, and attackers cannot access the complete training data. We refer to this scenario as data-constrained backdoor attacks. In such cases, previous attack methods suffer from severe efficiency degradation due to the entanglement between benign and poisoning features during the backdoor injection process. To tackle this problem, we introduce three CLIP-based technologies from two distinct streams: Clean Feature Suppression and Poisoning Feature Augmentation. The results demonstrate remarkable improvements, with some settings achieving over 100% improvement compared to existing attacks in data-constrained scenarios. Code is available at Data-constrained backdoor attacks.
11footnotetext: The first two authors contributed equally to this paper.22footnotetext: Corresponding Author: Bin Li.
[…]
@@ 1 Introduction
Deep neural networks (DNNs) are powerful ML algorithms inspired by the human brain, used in various applications such as image recognition He et al. (2016), natural language processing Liu et al. (2023), image generation Li et al. (2023b; 2022c; 2022b); Wu et al. (2023), and trajectory prediction Wong et al. (2022); Xia et al. (2022a). DNN effectiveness depends on training data quantity/quality. For example, Stable Diffusion (983M parameters) Rombach et al. (2022) excels in image generation due to pre-training on 5B image-text pairs. As the demand for data continues to rise, many users and businesses resort to third-party sources or online collections as a convenient means of acquiring the necessary data. However, recent studies Pan et al. (2022); Li et al. (2021b); Yang et al. (2021) have demonstrated that such practices can be maliciously exploited by attackers to contaminate the training data, significantly impairing the functionality and reliability of trained models.
The growing adoption of neural networks across different domains has made them an attractive target for malicious attacks. One particular attack technique gaining attention is the backdoor attack Goldblum et al. (2022); Nguyen & Tran (2020); Xia et al. (2022c). In backdoor attacks, a neural network is deliberately injected with a hidden trigger by introducing a small number of poisoning samples into the benign training set during the training. Once the model is deployed, the attacker can activate the backdoor by providing specific inputs containing the hidden trigger, causing the model to produce incorrect results. Backdoor attacks continue to present a significant and pervasive threat across multiple sectors, including image classification Gu et al. (2019), natural language processing Pan et al. (2022); Zeng et al. (2023), and malware detection Li et al. (2021a). In this paper, we focus on the widely studied field of image classification.
@@
It’s worth noting that previous backdoor attacks rely on a potentially overly broad assumption. They assume that all training data comes from a single source, and the collected source has been poisoned by attacker (as shown in O2O data collection mode in Fig. 1). This assumption grants attackers full access to the entire training dataset, making it easy to poison. However, it doesn’t accurately represent real-world attack scenarios. Consider a scenario where victims have a limited sample private dataset. To compensate, they may augment it by collecting additional data from various online sources (referred to as the public dataset) and combine it with their private data for training, as depicted in the M2O data collection mode in Fig. 1. In this case, some of the sources may be secretly poisoned by attackers. Attackers cannot access the private dataset and can only manipulate a portion of the public dataset for poisoning. Consequently, a discrepancy arises between the distribution of the poisoned data and the training data, deviating from previous poisoning attack pipeline.
@@
In this paper, we address a more realistic backdoor attack scenario called data-constrained backdoor attacks, where the attackers do not have access to the entire training set. To be more precise, we classify data-constrained backdoor attacks into three types based on different types of data sources: number-constrained backdoor attacks, class-constrained backdoor attacks, and domain-constrained backdoor attacks. Upon investigation, we have discovered that existing attack methods exhibit significant performance degradation when dealing with these data-constrained backdoor attacks. We propose that the entanglement between benign and poisoning features is a crucial factor contributing to this phenomenon. Entanglement refers to the neural networks utilizing both benign and poisoning features to make decisions for poisoning samples. However, this approach is not efficient for backdoor attacks. Ideally, an efficient backdoor attack should solely rely on the poison feature generated by the trigger to make decisions, irrespective of how the benign feature is expressed. To enhance the efficiency of poisoning attacks in data-constrained backdoor scenarios, we introduce two streams: Clean Feature Suppression and Poisoning Feature Augmentation to reduce the influence of clean features and amplify the expression of poisoning features, respectively. To achieve these goals, we propose three techniques utilizing the CLIP Radford et al. (2021).
@@
Our main contributions are summarized as follows. i) We present a novel and contemporary backdoor attack scenario called data-constrained backdoor attacks, which assumes that attackers lack access to the entire training data, making it a versatile and practical attack with broad applicability. ii) Through a systematic analysis of previous attack methods, we identify the entanglement between poisoning and benign features as the primary contributing factor to their performance degradation. iii) To address this issue, we introduce the pre-trained CLIP model into the field of backdoor attacks for the first time. We propose three innovative technologies: CLIP-CFE, CLIP-UAP, and CLIP-CFA. Extensive evaluations conducted on 3 datasets and 3 target models, and over 15 different settings demonstrate the significant superiority of our proposed CLIP-UAP and CLIP-CFA over existing backdoor attacks. Furthermore, CLIP-CFE complements existing attack methods and can be seamlessly integrated with them, resulting in further efficiency improvements.
2 Background and Related Work
Here, we provide a concise overview of the typical pipeline for backdoor attacks on neural networks. Sec. A.1 provides a comprehensive exploration of related work on backdoor attacks and CLIP.
@@ 2.1 Backdoor Attacks on Neural Networks
Backdoor attacks aim to introduce hidden triggers into DNNs, allowing the attacked models to behave correctly on clean samples while exhibiting malicious behavior when triggered by specific inputs. These attacks can occur at various stages of Artificial Intelligence (AI) system development Gao et al. (2020). The surface of backdoor attacks has been systematically categorized into six groups: code-based Bagdasaryan & Shmatikov (2021), outsourcing, pretrained model-based Wang et al. (2020); Ge et al. (2021), poisoning-based Liao et al. (2018), collaborative learning-based Nguyen et al. (2020), and post-deployment attacks Rakin et al. (2020). Among these categories, poisoning-based backdoor attacks, which involve introducing a backdoor trigger during the training process by mixing a few poisoning samples, are the most straightforward and commonly used method. This study focuses on addressing concerns related to poisoning-based backdoor attacks.
2.2 General Pipeline of Backdoor Attacks
@@
Backdoor injection. In this step, The attackers mix the poisoning set $\mathcal{P}$ into the clean training set $\mathcal{D}$ and release the new dataset. The victims download the poisoning dataset and use it to train their own DNN models Gu et al. (2019):
@@
Backdoor activation. In this step, the victims deploy their compromised DNN models on model-sharing platforms and model-selling platforms. The compromised model behaves normally when presented with benign inputs, but attackers can manipulate its predictions to align with their malicious objectives by providing specific samples containing pre-defined triggers.
@@ 3 Data-constrained Backdoor Attacks
We first show the considered pipeline of data-constrained backdoor attacks, and than illustrate the performance degradation of previous methods on proposed pipeline. Finally, we attribute the degradation to the entanglement between the benign and poisoning features during the poisoning injection.
@@ 3.1 Preliminaries
Previous methods Chen et al. (2017); Liu et al. (2017); Li et al. (2022a); Nguyen & Tran (2020) have commonly followed the attack pipeline outlined in Sec. 2.2. However, this widely adopted pipeline relies on an overly loose assumption that all training data is collected from a single source and that the attacker has access to the entire training data, which is often not the case in real-world attack scenarios. In this paper, we focus on a more realistic scenario: Data-constrained Backdoor Attacks.
@@
Pipeline of data-constrained backdoor attacks. The proposed pipeline also consists of three steps: poisoning set generation, backdoor injection, and backdoor activation. The backdoor injection and activation steps remain unchanged from the previous attack pipeline. However, in the poisoning set generation step, data-constrained attacks only assume access to a clean training set $\mathcal{D^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,N^{\prime}\}$, which follows a different data distribution from $\mathcal{D}$. To address this, the attacker randomly selects a subset $\mathcal{P^{\prime}}=\{(x_{i},y_{i})|i=1,\cdots,P\}$ from the accessible dataset $\mathcal{D^{\prime}}$, and creates the corresponding poisoning set $\mathcal{P}=\{(x^{\prime}_{i},k)|x^{\prime}_{i}=\mathcal{T}(x_{i},t),(x_{i},y_{i})\in\mathcal{P^{{}^{\prime}}},i=1,\cdots,P\}$. Additionally, based on the different constraints imposed by the accessible training set $\mathcal{D^{\prime}}$, data-constrained backdoor attacks are further categorized into three types: Number-constrained Backdoor Attacks, Class-constrained Backdoor Attacks, and Domain-constrained Backdoor Attacks. The details pipeline can be found in Sec. A.2.3, Sec A.2.4, and Sec. A.2.5, respectively.
@@
Empirical results. Noticing that the experimental setting of this experiment can be found in Sec. A.2.2. Fig. 2 (a), (b), and (c) illustrate the attack success rate on number-constrained, class-constrained, and domain-constrained backdoor attacks, respectively. The results demonstrate that ASR experiences a significant decrease as the number of poisoning samples ($P$), the number of class ($C^{\prime}$), or domain rate (the proportion of poisoning sets sampled from $\mathcal{D}\setminus\mathcal{D^{\prime}}$ and $\mathcal{D^{\prime}}$) in the poisoning set decreases, particularly for Blended backdoor attacks. It is worth noting that Universal Adversarial Perturbations (UAP) achieves relatively favorable results even with a low poisoning rate. This can be attributed to the utilization of a proxy model that is pre-trained on the entire training set ($\mathcal{D}$). However, UAP is not accessible in our settings, and we present the results for UAP to effectively demonstrate the performance degradation even when a pre-trained proxy model is available.
(a) Number-constrained
(b) Class-constrained
(c) Domain-constrained
@@
Figure 2: Attack success rate (ASR) in the different data-constrained backdoor attack. The experiment is repeated 5 times, and the solid lines represent the mean results. (a): The abscissa is the number ($P$) of samples in the poisoning set $\mathcal{P^{\prime}}$. (b): The experiments are conducted with triggers BadNets, Blended, and UAP, with a poisoning rate of $2\%$ ($P=1000$) for each. The x-axis represents the number of classes ($|Y^{\prime}|$) in the poisoning set $\mathcal{P^{\prime}}$. Specifically, ’1 (1)’ and ’1 (0)’ denote dirty-label single-class ($Y^{\prime}=\{c\},c\neq k$) and clean-label single-class ($Y^{\prime}=\{k\}$), respectively. (c): The poisoning rates of experiments with trigger BadNets, Blended, and UAP are $2\%$ ($P=1000$), $2\%$ ($P=1000$), and $1\%$ ($P=500$), respectively. The abscissa is the domain rate that represents the proportion of poisoning sets sampled from $\mathcal{D}\setminus\mathcal{D^{\prime}}$ and $\mathcal{D^{\prime}}$).